OpenStudy (anonymous):

Solve the system of equations. 2x + 3y = 1 y = 3x + 15 ( , )

OpenStudy (anonymous):

I will medal if its right

OpenStudy (anonymous):

Mathway.

OpenStudy (lalipapa123):

given, 2x+3y=1.......(1) y=3x+15 substitute y value in equation .1 2x+3(3x+15)=1 =>2x+9x+30=1 => 11x+30-1=0 11x+29=0 11x=29 x=29/11

OpenStudy (anonymous):

Where did you get 15? Im not saying that your wrong or anything I just want to understand how to do it

OpenStudy (anonymous):

Never mind Im an idiot I just realized the question had 15 in it

OpenStudy (lalipapa123):

hmm ! what yaar?

OpenStudy (anonymous):

Thank you

OpenStudy (lalipapa123):

confused me .am i right i.e my answer

OpenStudy (anonymous):

Wait how does 2 times (29) plus 3 times (11) = 1

OpenStudy (anonymous):

are they negative?